OpenStudy (anonymous):

how many solutions does the equation have? 2x+3y=1 -3x-2y=-1

OpenStudy (ihelpyoulearn):

Are there two equations? It looks like it but you said does the equation have?

OpenStudy (ihelpyoulearn):

@Justmeneedinghelp ^^^

OpenStudy (jango_in_dtown):

@Justmeneedinghelp This is a system of two linear equations in 2 unknown also the determinant value of the coefficient matrix is not zero. so the solution is unique... the solution is x=1/5, y=1/5

OpenStudy (irishboy123):

look at the equations. they look uncannily similar😋